# TiRE-GAN: Task-Incentivized Generative Learning Models for Radiomap Estimation with Radio Propagation Model

### v1

@@ Abstract
Enriching geometric information on radio frequency (RF) signal power distribution in wireless communication systems, the radiomap has become an essential tool for resource allocation and network management. Usually, a dense radiomap is reconstructed from sparse observations collected by deployed sensors or mobile devices, which makes the radiomap estimation an urgent challenge. To leverage both physical principles of radio propagation models and data statistics from sparse observations, this work introduces a novel task-incentivized generative learning model, namely TiRE-GAN, for radiomap estimation. Specifically, we first introduce a radio depth map as input to capture the overall pattern of radio propagation and shadowing effects, following which a task-driven incentive network is proposed to provide feedback for radiomap compensation depending on downstream tasks. Our experimental results demonstrate the power of the radio depth map to capture radio propagation information, together with the efficiency of the proposed TiRE-GAN for radiomap estimation.
@@
publicationid: pubid: 978-1-6654-3540-6/22 © 2022 IEEE
@@ I Introduction
The expansion of sensor networks and edge computing systems has stimulated many new technologies in the next generation of wireless communications, such as vehicle-to-everything (v2x) communications [1], unmanned aerial vehicle (UAV)-assisted networking [2] and fifth-generation (5G) terrestrial and non-terrestrial network (NTN) services [3]. All these applications require an accurate description of radio frequency (RF) spectrum coverage and efficient assessment of the radio environment, which lead to the new concept of radiomaps [4]. Describing the RF power spectral density (PSD) over different geometric locations, frequencies, and time, radiomaps can enrich the information of spectrum coverage and assist the aforementioned spectrum management applications [5]. Usually, dense radiomaps are reconstructed from sparse observations collected by mobile sensors and user devices, as shown in Fig. 1. To fully leverage the power of radiomaps in spectrum resource allocation and network optimization, efficient radiomap estimation (RME) from sparse observations has emerged as an urgent challenge.
@@
Existing methods of radiomap estimation can be categorized into model-based or learning-based approaches, each bearing its distinct constraints. Model-based methods usually assume a specific radio propagation model, with which unknown parameters are estimated by minimizing the error between observed samples and predictions. Typical examples include the log-distance path loss (LDPL) models [6], inverse distance weighted interpolation [7] and thin-plate splines kernels [8]. However, the performances of model-based RME highly rely on the selection of physical radio propagation models, which are usually unknown in practical applications. Moreover, the PSD distribution sometimes is more sensitive to surrounding obstacles rather than following a strict empirical model, especially in complicated environments. The efficient extraction of radio model information from realistic data samples remains an open question.
@@
With the development of artificial intelligence (AI), recent interests of RME focus on learning-based methods [9]. In [10], a fast radiomap estimation is proposed based on Unet. Another work utilizing deep convolutional neural networks is [11], where an auto-encoder structure is presented for radiomap estimation. Other learning-based methods also include the conditional generative adversarial nets (cGAN) [3] and transformers [12]. Despite many successes, efficient learning-based RME requires sufficient data samples with satisfying quality, which are usually inaccessible in realistic scenarios. Although some works [13, 14] consider integrating the radio propagation models with learning machines to address the data scarcity, the efficient model selection remains challenging. Moreover, how to satisfy the special requirements of downstream tasks, such as outage detection and network planning, is another critical concern.
@@
To address the aforementioned problems, this work proposed a Task-incentivized Radiomap Estimation based on Generative Adversarial Networks (TiRE-GAN) for radiomap estimation, leveraging the synergies of radio propagation model and generative AI. Specifically, a radio depth map is introduced as an additional input channel of the generator to capture the overall radio propagation information and shadowing effects. Then, a task-incentivized block is proposed to provide feedback for radiomap generation to compensate for the estimated PSD based on downstream tasks. Our contributions can be summarized as follows:
  • We introduce a novel radio depth map (RDM) as input conditions to the cGAN, serving as a proxy for the shadowing effects and radio propagation behavior, thereby infusing the learning process with additional insights into the physical principles of radio propagation.
  • We integrate the physics-conditioned cGAN with a novel task-incentivized network to provide compensation to the estimated radiomaps depending on downstream tasks, which significantly enhances the RME performance.
  • Our experimental results demonstrate the power of the proposed radio depth map in characterizing radio propagation behavior, as well as the efficiency of the proposed TiRE-GAN in radiomap estimation.
@@ II Problem Description
Figure 1: Problem Description of RME: Given the distribution of buildings, transmitter locations (in read), and the sparse observations collected by mobile sensors and devices, we aim to predict the dense radiomaps.
Figure 2: Overall Structure of TiRE-GAN consisting of 3 parts: 1) The model-based data preprocessing module utilizes the input observed information (e.i. building map and transmitter location) to generate the radio depth map and then stacks the four layers together as input; 2) The radiomap (RM) generation module is in a cGAN generator structure, which learn and produce the prediction of radiomap; 3) The feedback module consists of 2 parts: a cGAN discriminator and a task-incentivized network. Both modules produce a loss term as feedback to improve the generator.
In this section, we first introduce our problem formulation. As shown in Fig. 1, the radiomap estimation is formulated as a sparse sensing problem, where a dense radiomap $\hat{M}_{R}$ is reconstructed given the landscape information of the surrounding environment, sparse observations collected from sensors and mobile devices, and locations of transmitters, formulated by the following equation:
@@
where $M_{U}$, $M_{T}$, $M_{S}$ and $\mathcal{X}$ represent the landscape map, the encoded transmitter location, the sampled radiomap, and additional extracted features, respectively.
@@
In this work, our objectives can be summarized as two parts: 1) design an effective mapping function $g(\cdot)$ by integrating the cGAN with a novel task-incentivized network; and 2) design a suitable loss function $L(\cdot)$ to embed the physical principles of radio propagation. We shall introduce the details of our proposed TiRE-GAN in Section III.
@@ III-A Overall Structure of TiRE-GAN Framework
We now introduce the overall structure of our proposed TiRE-GAN. As shown in Fig. 2, the architecture of the proposed framework can be structured into three components: 1) model-based data preprocessing; 2) radiomap generation based on cGAN; and 3) task-incentivized feedback.
The overall operational sequence of the training process is delineated as follows. To embed the radio propagation model, the urban landscape map and the transmitter location data are processed to fabricate a physics-embedded Radio Depth Map (RDM), which will be further discussed in Section III-B2. Subsequently, this RDM, alongside the sampled radiomap, urban map, and transmitter location data, are amalgamated into a quadruple-channel input for the cGAN. Throughout the training phase, a pre-trained task-incentivized network is incorporated as a contributory factor within the loss function for the cGAN’s generator, enhancing the model’s ability to compensate for the radiomap estimations by embedding physics-based constraints and the rewards from downstream tasks. We will then introduce each module as follows.
@@ III-B1 Observed Data Preprocessing
Suppose that the radiomap is formatted in a $N\times N$ grid, i.e., $M_{R}\in\mathbb{R}^{N\times N}$. The landscape map of buildings is encoded as a binary segmented image $M_{U}\in\mathbb{R}^{N\times N}$, where building areas are marked by 1 and open areas are annotated by 0. Similar to the building map, we encode the transmitter locations into a binary image $M_{T}\in\mathbb{R}^{N\times N}$, where the transmitter locations are marked as 1; otherwise, the pixel values are zero. The sparse observations can be viewed as the corresponding pixels in gray-scale images $M_{S}\in\mathbb{R}^{N\times N}$. For convenience, we first use a threshold to eliminate the extremely low-value observations, and then normalize the PSD value into $0\sim 1$ for the observed locations. Since linear normalization is applied, we can easily re-scale the estimated radiomap back to the original data domain. For the missing values in $M_{S}$, we set them as zeros. Since a convolutional neural network is applied for the cGAN structure, zero-padding does not impact the final radiomap estimation.
@@ III-B2 Radio Depth Map
To embed the radio propagation information, we introduce the Radio Depth Map (RDM), $\mathcal{X}=M_{D}\in\mathbb{R}^{N\times N}$ as a novel model-based attribute, derived from the distribution of buildings and the spatial locations of transmitters. Similar to a depth map in computer vision which provides information on the distances of surfaces from a specific viewpoint, the RDM is formulated as:
$M_{D}(x,y)=n\left(\sum_{t=1}^{N_{T}}P_{t}(x,y)\cdot B_{t}(x,y)\right),$ (3)
where $(x,y)$ denotes the two-dimensional coordinates of the target location, $P_{t}(x,y)$ denotes the distance-based path loss gain from the $t$-th transmitter, and the $B_{t}(x,y)$ capture the impact of buildings from the viewpoint of $t$-th transmitter.
Suppose that $C(l,l_{t})$ identifies the set of pixels in the direct line of sight between the target location $l=(x,y)$ and location $l_{t}=(x_{t},y_{t})$ of the $t$-th transmitter. The shadowing term impacted by buildings is calculated by
$B_{t}(x,y)=\frac{\sum_{i\in C(l,l_{t})}(1-M_{U}(i))}{\sum_{i\in C(l,l_{t})}1},$ (4)
which characterizes the ratio of non-building pixels between $l$ and $l_{t}$. $B_{t}(x,y)$ is larger and the shadowing effect is smaller if the radio propagates through fewer buildings.
To capture the path loss $P_{t}(x,y)$, we employ the Inverse Distance Weighting (IDW) model, calculated by
$P_{t}(x,y)=d(l,l_{t})^{-\lambda},\ \lambda>0,$ (5)
where $d(l,l_{t})$ denotes the Euclidean distance between the target point $l$ and the $t$-th the transmitter. The parameter $\lambda$ governs the rate, at which signal strength diminishes with increasing distance, necessitating a positive value to ensure physical plausibility in the depth map. An exemplary radio depth map is shown in the orange block of Fig. 2.
@@
In this work, we employ a conditional Generative Adversarial Network (cGAN) [15] as the foundational architecture, which consists of two principal components: a generator $G$ and a discriminator $D$. Different from the traditional GAN which generates data from random noise, the generator in cGAN can generate data from given features $\mathcal{F}$ and incorporate additional information $\mathbf{I}$ in the discriminator. In our radiomap estimation, the conditional features consist of four channels as processed in the Data Preprocessing Module, i.e., =$\mathcal{F}=\{M_{U},M_{T},M_{S},M_{D}\}$ as shown in Fig. 2. The additional information $\mathbf{I}$ can be designed as a function of the region features, i.e., $\mathbf{I}=h(\mathcal{F})$ or some determined labels, such as fake or real [16]. Then, the basic cGAN can be formulated as
@@
where $p_{\rm data}(\bf{y,I})$ is the joint distribution with the prior knowledge $\mathbf{I}$, and $\mathbf{y}$ is the generated data. Generally, the optimization of a min-max problem in cGAN can be split into two sub-problems, and train $D$ and $G$ iteratively. Specifically, the loss function to optimize $D$ is
$L_{D}=-V(D,G),$ (7)
and the general loss function to optimize $G$ is
@@
Here, $L_{MSE}$ describes the error of RME by Mean Squared Error (MSE), i.e.,
$L_{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}(y_{i}-\hat{y}_{i})^{2},$ (10)
where $K$ is the total number of pixels, $\mathbf{y}$ and $\hat{\mathbf{y}}$ are the ground truth and the predicted radiomap, while $y_{i}$ and $\hat{y}_{i}$ are the true and predicted pixel value of the $i$th pixel. $L_{R}$ is a task-incentivized regularization term from the feedback module, which will be introduced in Section III-D.
@@ III-D Feedback Design
$\vspace{-5mm}L_{D}(\mathbf{y},\hat{\mathbf{y}})=-\frac{1}{K}\sum_{i=1}^{K}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right].$ (11)
Figure 3: Task-incentivized network (TIN) structure: We design the TIN based on Unet, which contains 4 encoding blocks and 4 decoding blocks. Each encoding block consists of 3 convolution layers and a pooling layer. Each decoding block consists of one concatenation layer, a transpose convolution layer, and 3 convolution layers.
@@ III-D2 Task-Incentivized Network
In traditional RME, the objective transcends the mere accuracy of radio signal strength at discrete locations within a dense radiomap. However, in practical scenarios, the final objective of RME is to assist spectrum planning and network management. For example, one can easily identify the outage of regions based on their radiomap by setting a certain threshold of service failure. To compensate for the radiomap estimations based on downstream tasks, we introduce a TIN module as shown in Fig. 2.
As an exemplary TIN, we adopt outage detection as a downstream task. More specifically, the proposed TIN utilizes a UNet architecture for semantic segmentation of radio power outage as Fig. 3, where the input constitutes a radiomap and the output is an outage map indicating signal coverage or absence. The training of the TIN is conducted using ground-truth radio coverage and outage maps, thereby ensuring that the network learns to accurately predict outage areas based on real-world data. Once trained, the TIN operates within the cGAN framework as a pre-trained component. During inference, the output generated by the cGAN’s generator serves as the input to the TIN, enabling it to contribute an additional term $L_{R}$ to the loss function of the cGAN. This term is specifically designed to encapsulate aspects of human-centric interest, such as the distribution and characteristics of signal outages, thereby guiding the cGAN toward generating radiomaps that are not only precise but also aligned with the real-world considerations and requirements of downstream tasks. This integration of the TIN within the cGAN architecture exemplifies a strategic convergence of machine learning precision with domain-specific insights, enhancing the utility and applicability of the generated radiomaps.
The TIN provides the task-incentivized regularization term $L_{R}$ in the loss function of the generator, which is calculated by the pre-trained TIN using the MSE loss to measure the error between the predicted and the ground truth outage maps:
$L_{R}(z,\hat{z})=\frac{1}{K}\sum_{i=1}^{K}(z_{i}-\hat{z}_{i})^{2},$ (12)
where the $z$ and $\hat{z}$ are the ground truth and the predicted outage map, while $z_{i}$ and $\hat{z}_{i}$ are the true and predicted pixel values of the $i$th pixel. $K$ is the number of pixels.
Note that, our TIN can be integrated with any RM-assisted applications. Some other typical examples also include ray tracing or wireless localization.
Figure 4: Performance comparison (sr=1) between Kriging, matrix completion(MC), AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), our TiRE-GAN, and the ground truth (GT) on 3 radiomap examples. TiRE-GAN achieves the best radio propagation and the shadowing effect prediction.

### v3

# TiRE-GAN: Task-Incentivized Generative Learning for Radiomap Estimation

@@ Abstract
To characterize radio frequency (RF) signal power distribution in wireless communication systems, the radiomap is a useful tool for resource allocation and network management. Usually, a dense radiomap is reconstructed from sparse observations collected by deployed sensors or mobile devices. To leverage both physical principles of radio propagation models and data statistics from sparse observations, this work introduces a novel task-incentivized generative learning model, namely TiRE-GAN, for radiomap estimation. Specifically, we first introduce a radio depth map to capture the overall pattern of radio propagation and shadowing effects, following which a task-driven incentive network is proposed to provide feedback for radiomap compensation depending on downstream tasks. Our experimental results demonstrate the power of the radio depth map to capture radio propagation information, and the efficiency of the proposed TiRE-GAN for radiomap estimation.
@@
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Rapid growth of sensor networks and edge computing continue to advance the development of next-generation wireless communications, stimulating many novel technologies, such as vehicle-to-everything [1] and integrated sensing and communications [2]. All these applications require an accurate description of radio frequency (RF) spectrum coverage and efficient assessment of the radio environment, leading to the concept of radiomaps [3]. Radiomaps describe RF power spectral density (PSD) across various locations, frequencies, and time, offering rich information of spectrum coverage for the spectrum management applications. Usually, dense radiomaps are reconstructed from sparse observations collected by mobile sensors. To fully leverage the power of radiomaps in spectrum management, efficient radiomap estimation (RME) from sparse observations has emerged as an urgent challenge.
@@
Existing methods of radiomap estimation can be categorized into model-based or learning-based approaches, each bearing its distinct constraints. Model-based approaches, such as the log-distance path loss (LDPL) models [4], inverse distance weighted interpolation (IDW) [5], and thin-plate splines kernels [6], usually assume a specific radio propagation model, with which unknown parameters are estimated by minimizing the error between observed samples and predictions. However, in practical applications, these physics models are often unknown, making model selection challenging. Moreover, the PSD distribution sometimes is more sensitive to surrounding obstacles rather than following a strict empirical model, especially in complicated environments. Thus, the efficient extraction of radio model information from realistic data samples remains an open question.
@@
With the development of artificial intelligence (AI), recent interests of RME focus on learning-based methods [7]. In [8], a fast radiomap estimation is proposed based on Unet. Another work utilizing deep convolutional neural networks is [9], where an auto-encoder structure is presented for radiomap estimation. Other learning-based methods also include conditional generative adversarial nets (cGAN) [10, 11] and transformers [12]. However, these learning-based methods require high-quality, abundant data, which is usually inaccessible in practical applications. Although some hybrid methods [13, 14] attempt to combine radio models with machine learning approaches, the efficient model selection and task-specific designed RME, such as for outage detection and network planning, remain critical concerns.
@@
To address the aforementioned problems, we propose a Task-incentivized Radiomap Estimation using Generative Adversarial Networks (TiRE-GAN), leveraging the synergies of radio propagation model and generative AI. Specifically, a radio depth map is introduced as an additional input channel of the generator to capture the overall radio propagation information and shadowing effects. Then, a task-incentivized block is proposed to provide feedback for radiomap generation to compensate for the estimated PSD based on downstream tasks. Our contributions can be summarized as follows:
We introduce a novel radio depth map (RDM) as an input to the cGAN, capturing shadowing and radio propagation effects to enhance the learning process with physical insights. We develop a physics-conditioned cGAN with a task-incentivized network, which improves RME performance by compensating radiomaps based on downstream tasks. Experiments demonstrate the effectiveness of the proposed RDM in capturing radio propagation behavior and the efficiency of TiRE-GAN.
@@ II Problem Description
Figure 1: Overall Structure of TiRE-GAN consisting of 4 parts: 1) The model-based data preprocessing module utilizes the input observed information (i.e. building map and transmitter location) to generate the radio depth map and then stacks the four layers together as input; 2) The radiomap (RM) generation module is in a cGAN generator structure, which learn and produce the prediction of radiomap; 3) The feedback module consists of 2 parts: a cGAN discriminator and a task-incentivized network. Both modules produce a loss term as feedback to improve the generator; 4) The Downstream task that provides the additional loss term for training guidance.
In this section, we first introduce our problem formulation. In this work, the radiomap estimation is formulated as a sparse sensing problem, where a dense radiomap $\hat{M}_{R}$ is reconstructed given the landscape information of the surrounding environment, sparse observations collected from sensors, and locations of transmitters, formulated by the following equation:
@@
where $g(\cdot)$ denotes the RME neural network. Here, $M_{U}$, $M_{T}$, and $M_{S}$ represent the landscape map, the encoded transmitter location, and the sampled radiomap, respectively. $\mathcal{X}=f_{physical}(M_{U},M_{T})$ is a function embedding the physical principles of radio propagation behavior. In this work, radio depth map is applied to capture the $f_{physcal}(\cdot)$, which will be introduced with details in Section III-B2.
@@
In this work, our objectives are to: 1) design an effective mapping function $g(\cdot)$ by integrating the cGAN with a novel task-incentivized network; and 2) design a suitable loss function $L(\cdot)$ to embed the physical principles of radio propagation. We introduce the details of our proposed TiRE-GAN in Section III.
@@ III-A Overall Structure of TiRE-GAN Framework
We now introduce the overall structure of our proposed TiRE-GAN. As shown in Fig. 1, the architecture of the proposed framework can be structured into three components: 1) model-based data preprocessing; 2) radiomap generation; and 3) task-incentivized feedback. The training process operates as follows. First, the urban landscape map and transmitter location data are processed to construct a physics-embedded Radio Depth Map (RDM), detailed further in Section III-B2, after which it is fed into the cGAN generator together with the urban map, transmitter locations and sparse observations. During training, a pre-trained task-incentivized network is integrated to provide feedback by embedding physics-based constraints and rewards from downstream tasks. Each module will be introduced as follows.
@@ III-B1 Observed Data Preprocessing
Suppose that the radiomap is formatted in a $N\times N$ grid, i.e., $M_{R}\in\mathbb{R}^{N\times N}$. The landscape map of buildings is encoded as a binary segmented image $M_{U}\in\mathbb{R}^{N\times N}$, where building areas are marked by 1 and open areas are annotated by 0. Similar to the building map, we encode the transmitter locations into a binary image $M_{T}\in\mathbb{R}^{N\times N}$, where the transmitter locations are marked as 1; otherwise, the pixel values are zero. The sparse observations can be viewed as a grayscale image $M_{S}\in\mathbb{R}^{N\times N}$ via zero-padding, where we use the observed PSD the as corresponding pixel values; otherwise, the pixel value is zero. For convenience, we first use a threshold to eliminate the extremely low-value observations, and then normalize the PSD value into $0\sim 1$ for the observed locations. Since linear normalization is applied, we can easily re-scale the estimated radiomap back to the original data domain. For the missing values in $M_{S}$, we set them as zeros. Since a convolutional neural network is applied for the cGAN structure, zero-padding does not impact the final RME.
@@ III-B2 Radio Depth Map
In Eq. (1), we included an $\mathcal{X}$ of additional extracted features for RME. To embed the radio propagation information, we introduce the Radio Depth Map (RDM), $\mathcal{X}=M_{D}\in\mathbb{R}^{N\times N}$ as a novel model-based attribute, derived from the distribution of buildings and the spatial locations of transmitters. Similar to a depth map in computer vision which provides information on the distances of surfaces from a specific viewpoint, the RDM is formulated as:
$M_{D}(\xi,\eta)=n\left(\sum_{t=1}^{N_{T}}P_{t}(\xi,\eta)\cdot B_{t}(\xi,\eta)\right),$ (3)
where $(\xi,\eta)$ denotes the two-dimensional coordinates of the target location, $N_{T}$ is the number of transmitters, $P_{t}(\xi,\eta)$ denotes the distance-based path loss gain from the $t$-th transmitter, $n(\cdot)$ is the max normalization function to re-scale the maximal value in the RDM as $1$, and the $B_{t}(\xi,\eta)$ capture the impact of buildings from the viewpoint of $t$-th transmitter.
Let $C(\ell,\ell_{t})$ denote the set of pixels in the direct line of sight between the target location $\ell=(\xi,\eta)$ and location $\ell_{t}=(\xi_{t},\eta_{t})$ of the $t$-th transmitter. The shadowing term impacted by buildings is calculated by
$B_{t}(\xi,\eta)=\frac{\sum_{i\in C(\ell,\ell_{t})}(1-M_{U}(i))}{\sum_{i\in C(\ell,\ell_{t})}1},\vspace{-1mm}$ (4)
which characterizes the ratio of non-building pixels between $\ell$ and $\ell_{t}$. $B_{t}(\xi,\eta)$ is larger and the shadowing effect is smaller if the radio propagates through fewer buildings.
To capture the path loss $P_{t}(\xi,\eta)$, we employ the Inverse Distance Weighting (IDW) model, calculated by
$P_{t}(\xi,\eta)=d(\ell,\ell_{t})^{-\lambda},\ \lambda>0,$ (5)
where $d(\ell,\ell_{t})$ denotes the Euclidean distance between the target $\ell$ and the $t$-th transmitter. The parameter $\lambda$ governs the rate, at which signal strength diminishes with increasing distance, necessitating a positive value to ensure physical plausibility in the depth map. An exemplary radio depth map is shown in the orange block of Fig. 1.
@@
In this work, we employ a cGAN [15] as the foundational architecture, which consists of two principal components: a generator $G$ and a discriminator $D$. The generator employs a ResNet-based architecture, consisting of 6 ResNet blocks. The discriminator uses a PatchGAN structure, comprising 4 convolutional blocks , and ends with a Sigmoid activation to classify the input as real or generated. In our radiomap estimation, the conditional features consist of four channels as processed in the model-based data preprocessing module, i.e., =$\mathcal{F}=\{M_{U},M_{T},M_{S},M_{D}\}$ as shown in Fig. 1. The additional information $\mathbf{I}$ can be designed as a function of the region features, i.e., $\mathbf{I}=h(\mathcal{F})$ or some determined labels, such as surrounding landscape information [16]. Then, the basic cGAN can be formulated as
@@
where $p_{\rm data}(\bf{y,I})$ is the joint distribution with the prior knowledge $\mathbf{I}$, and $\mathbf{y}$ is the true radiomap. $p_{\rm data}(\mathbf{\mathcal{F}})$ denotes the distribution of input features used for cGAN training. Interested readers could refer to [16, 10] for a detailed introduction to cGAN structure. Generally, the optimization of a min-max problem in cGAN can be split into two sub-problems, and train $D$ and $G$ iteratively. Specifically, the loss function to optimize $D$ is $L_{D}=-V(D,G)$, and the general loss function to optimize $G$ is
@@
Here, $L_{MSE}$ describes the error of RME by Mean Squared Error (MSE), i.e., $L_{MSE}(\mathbf{y},\hat{\mathbf{y}})=\frac{1}{K}\sum_{i=1}^{K}(y_{i}-\hat{y}_{i})^{2}$, where $K$ is the total number of pixels, $\mathbf{y}$ and $\hat{\mathbf{y}}$ are the ground truth and the predicted radiomap, while $y_{i}$ and $\hat{y}_{i}$ are the true and predicted pixel value of the $i$th pixel. $L_{R}$ is a task-incentivized regularization term from the feedback module, which will be introduced in Section III-D.
@@ III-D Feedback Design
$L_{D}(\mathbf{\delta},\hat{\mathbf{\delta}})=-\frac{1}{K}\sum_{i=1}^{K}\left[\delta_{i}\log(\hat{\delta}_{i})+(1-\delta_{i})\log(1-\hat{\delta}_{i})\right],\vspace{-4mm}$ (9)
where $\delta$ and $\hat{\delta}$ are the true and predicted values. $K$ is the number of data points in the batch.
Figure 2: Task-incentivized network (TIN): based on Unet, TIN contains 4 encoding blocks and 4 decoding blocks. Each encoding block uses 3 convolution layers and a pooling layer. Each decoding block consists of one concatenation layer, a transpose convolution layer, and 3 convolution layers.
@@ III-D2 Task-Incentivized Network
The goal of TiRE-GAN extends beyond capturing radio signal strength at specific locations; it shall ultimately support spectrum planning and network management. For instance, radiomaps can highlight service outages by setting a failure threshold. To enhance radiomap estimations for downstream tasks, we introduce a TIN module, as shown in Fig. 1.
We use outage detection as an exemplary TIN, employing a UNet architecture for semantic segmentation of radio power outages as shown in Fig. 2. The input is a radiomap, and the output outage map indicates signal coverage. Trained on real radio coverage and outage maps, the TIN accurately predicts outage areas. The TIN operates within the cGAN framework as a pre-trained component, for which the generator’s output serves as input. The TIN then contributes an additional loss term, $L_{R}$, guiding cGAN to produce radiomaps that not only maintain precision but also address real-world concerns relevant to downstream tasks.
The TIN provides a task-incentivized regularization term $L_{R}$ in the loss function of the generator, which is calculated by the pre-trained TIN using the MSE loss to measure the error between the predicted and the ground truth outage maps, i.e.,
$L_{R}(z,\hat{z})=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}(z_{i}-\hat{z}_{i})^{2},\vspace{-2mm}$ (10)
where the $z$ and $\hat{z}$ are the ground truth and the predicted outage map, while $z_{i}$ and $\hat{z}_{i}$ are the true and predicted pixel values of the $i$th pixel. $N_{p}$ is the number of pixels.
Note that our TIN can be integrated with any RM-assisted applications. Some typical examples include ray tracing or wireless localization.
Figure 3: Performance comparison (sr=1) between Kriging, AutoEncoder (AE), RadioUNet (RU), conditional GAN (cGAN), RME-GAN, Geo2SigMap, our TiRE-GAN, and the ground truth (GT) on 3 radiomap examples. TiRE-GAN achieves the best radio propagation and the shadowing effect prediction.